# Laplace Transforms of Stopping Times for Subordinator with Applications to Inventory Control

### v1

@@ Abstract.
Intermittent demand fluctuations pose significant challenges in disaster logistics and medical supply systems. In this study, we formulate cumulative demand as a generalized Lévy process composed of a drift term, Poisson jumps, and compound Poisson jumps, and analyze a continuous-time inventory model. The proposed framework provides a unified formulation that encompasses both drifted Poisson processes and drifted compound Poisson processes.
From a mathematical perspective, we treat the reorder time as a first-passage problem of a subordinator and derive its Laplace transform via the Laplace exponent. In particular, for the drifted Poisson case, we obtain an explicit representation of the inverse Laplace exponent using the Lambert W function, which yields an analytic expression for the Laplace transform of the first-passage time. Furthermore, when the jump sizes follow exponential and Gamma distributions, we derive explicit formulas for the mean and variance of the reorder times, thereby clarifying the moment structure of first-passage times for generalized Lévy demand processes.
From an operations research perspective, we explicitly characterize the expected total cost over a finite time horizon based on the distribution of cumulative demand. This study presents an analytical framework that integrates first-passage theory of Lévy processes with continuous-time inventory control.
@@ 1. Introduction
Inventory control problems under uncertain demand constitute an important research area in both probability theory and operations research. In particular, in situations such as disaster response or intermittent demand, continuous demand models are often insufficient. Lévy processes provide a well-behaved class of stochastic processes capable of modeling discontinuous changes.
@@ 2.1. Fixed-Order-Quantity Policy under Intermittent Demand
Intermittent demand refers to a demand pattern in which demand does not occur continuously, but instead exhibits sudden arrivals after long periods of no demand. Such demand is characterized by strong uncertainty in both the timing of demand occurrences and the demand sizes. We model the cumulative demand under intermittent demand by a subordinator $\{D_{t}\}_{t\geq 0}$.
In this model, the inventory level at time $t$ is given by the initial inventory minus the cumulative demand up to time $t$. Whenever the inventory level falls below a prescribed threshold, a replenishment of fixed size $Q$ is placed. Since the time at which the inventory reaches the threshold is stochastic, we define the corresponding stopping times.
To model replenishment at each reorder point, the system is designed so that an order is placed whenever the cumulative demand exceeds a prescribed level. The order quantity is also incorporated explicitly. The formulation follows [5].
@@ 3.1. Case where the Cumulative Demand is a Drifted Poisson Process
$E[T_{n}]$ $=\frac{a+(n-1)Q}{\mu+\alpha\lambda},$
$\mathrm{Var}[T_{n}]$ $=\frac{\alpha^{2}\lambda(a+(n-1)Q)}{\left(\mu+\alpha\lambda\right)^{3}}.$
@@
be defined.
@@
Let $B=\frac{\alpha}{\mu}$ and $C=\frac{\alpha\lambda}{\mu}$.
@@ 3.2. Case where the Cumulative Demand is a Drifted Compound Poisson Process
be defined.
@@
$\quad\frac{d^{2}\Phi}{ds^{2}}(s)$
$=-\frac{1}{2\mu}\times\frac{\left\{\sqrt{(\mu\eta+s+\lambda)^{2}-4\mu s\eta}-(s+\lambda-\mu\eta)\frac{(s+\lambda-\mu\eta)}{\sqrt{(\mu\eta+s+\lambda)^{2}-4\mu s\eta}}\right\}}{\left\{(\mu\eta+s+\lambda)^{2}-4\mu s\eta\right\}^{\frac{3}{2}}}$
@@
Let $L_{J_{1}}=\left(\frac{\eta}{\eta-\theta}\right)^{\beta}$. Then, for $\theta<\eta$,
@@
Let $\psi^{-1}=\Phi$ denote the inverse function of $\psi$. Then,
$\Phi(0)=0\quad(\because\psi(0)=0).$
$\frac{d\Phi}{ds}(0)=\frac{1}{\cfrac{d\psi}{d\theta}(0)}=\frac{1}{\mu+\lambda\beta\eta^{-1}}=\frac{\eta}{\mu\eta+\beta\lambda}.$
@@
Therefore,
$E[T_{n}]$ $=-\left.\frac{d}{ds}E[e^{-sT_{n}}]\right|_{s=0}$
$=K\exp(-K\Phi(0))\frac{d\Phi}{ds}(0)$
$=\frac{\eta(a+(n-1)Q)}{\mu\eta+\beta\lambda}.$
@@ 3.3. Generalization of the Cumulative Demand
Let $\mu,\alpha,\lambda,\lambda^{\prime}>0$. Let $\{N_{t}\}_{t\geq 0}$ be a $\lambda$-Poisson process, $\{N^{\prime}_{t}\}_{t\geq 0}$ be a $\lambda^{\prime}$-Poisson process, and $\{J_{k}\}_{k\geq 1}$ be a sequence of independent and identically distributed random variables. Assume that $\{N_{t}\}_{t\geq 0}$, $\{N^{\prime}_{t}\}_{t\geq 0}$, and $\{J_{k}\}_{k\geq 1}$ are mutually independent.
We define the cumulative demand process $\{D_{t}\}_{t\geq 0}$ as
@@
we derive $E[e^{-sT_{n}}]$, $E[T_{n}]$, and $\text{Var}[T_{n}]$.
Let $\psi(\theta)$ denote the Laplace exponent of $D_{t}$. Then,
@@
Since $L_{J_{1}}=\left(\frac{\eta}{\eta-\theta}\right)^{\beta}$, for $\theta<\eta$,
$\psi(\theta)=\log E[e^{\theta D_{1}}]=\theta\mu+\lambda(e^{\theta\alpha}-1)+\lambda^{\prime}\left(\left(\frac{\eta}{\eta-\theta}\right)^{\beta}-1\right).$
$\frac{d\psi}{d\theta}(\theta)=\mu+\alpha\lambda e^{\alpha\theta}+\lambda^{\prime}\beta\eta^{\beta}(\eta-\theta)^{-(\beta+1)}.$
@@
Let $\Phi$ denote the inverse function of $\psi$. Then,
$\Phi(0)=0\quad(\because\psi(0)=0).$
$\frac{d\Phi}{ds}(0)=\frac{1}{\cfrac{d\psi}{d\theta}(0)}=\frac{1}{\mu+\alpha\lambda+\lambda^{\prime}\beta\eta^{-1}}=\frac{\eta}{\mu\eta+\alpha\lambda\eta+\beta\lambda^{\prime}}.$
@@
Therefore,
$E[T_{n}]$ $=-\left.\frac{d}{ds}E[e^{-sT_{n}}]\right|_{s=0}$
$=K\exp(-K\Phi(0))\frac{d\Phi}{ds}(0)$
$=\frac{\eta(a+(n-1)Q)}{\mu\eta+\alpha\lambda+\beta\lambda^{\prime}}.$
@@ 3.4. On the Expected Total Cost
Although we do not directly use the distribution of $T_{n}$, we consider the generalized cumulative demand process and investigate the expected value of the total inventory cost derived from it.
Consider the cumulative demand process
@@
follows a Gamma distribution with parameters $j\beta,\eta$.
Hence,
@@
Here, $\{N_{t}\}_{t\geq 0}$ is a $\lambda$-Poisson process, $\{N^{\prime}_{t}\}_{t\geq 0}$ is a $\lambda^{\prime}$-Poisson process, and $\{J_{k}\}_{k\geq 1}$ is a sequence of nonnegative independent and identically distributed random variables. Assume that these are mutually independent.
Then, for any $s>0$ and $b>0$,
@@
holds.
In particular, when $\beta=1$,
@@
Theorem 3.17 ((K.) Comparison between Finite-Horizon and Long-Run Average).

### v2

# arXiv:2602.19625v2[math.PR] 03 Mar 2026

@@ Abstract.
Intermittent demand fluctuations pose significant challenges in disaster logistics and medical supply systems. In this study, we formulate cumulative demand as a generalized Lévy process composed of a drift term, Poisson jumps, and compound Poisson jumps, and analyze a continuous-time inventory model. The proposed framework provides a unified formulation that encompasses both drifted Poisson processes and drifted compound Poisson processes. From a mathematical perspective, we treat the reorder time as a first-passage problem of a subordinator and derive its Laplace transform via the Laplace exponent. In particular, for the drifted Poisson case, we obtain an explicit representation of the inverse Laplace exponent using the Lambert W function, which yields an analytic expression for the Laplace transform of the first-passage time. Furthermore, when the jump sizes follow exponential and Gamma distributions, we derive explicit formulas for the mean and variance of the reorder times, thereby clarifying the moment structure of first-passage times for generalized Lévy demand processes. From an operations research perspective, we explicitly characterize the expected total cost over a finite time horizon based on the distribution of cumulative demand. This study presents an analytical framework that integrates first-passage theory of Lévy processes with continuous-time inventory control.
@@ 1. Introduction
Inventory control problems under uncertain demand constitute an important research area in both probability theory and operations research. In particular, in situations such as disaster response or intermittent demand, continuous demand models are often insufficient. Lévy processes provide a well-behaved class of stochastic processes capable of modeling discontinuous changes. ( [9][6][10][5][4])
@@ 2.1. Fixed-Order-Quantity Policy under Intermittent Demand
Intermittent demand refers to a demand pattern in which demand does not occur continuously, but instead exhibits sudden arrivals after long periods of no demand. Such demand is characterized by strong uncertainty in both the timing of demand occurrences and the demand sizes. We model the cumulative demand under intermittent demand by a subordinator $\{D_{t}\}_{t\geq 0}$. In this model, the inventory level at time $t$ is given by the initial inventory minus the cumulative demand up to time $t$. Whenever the inventory level falls below a prescribed threshold, a replenishment of fixed size $Q$ is placed. Since the time at which the inventory reaches the threshold is stochastic, we define the corresponding stopping times. To model replenishment at each reorder point, the system is designed so that an order is placed whenever the cumulative demand exceeds a prescribed level. The order quantity is also incorporated explicitly. The formulation follows [5].
@@ 2.2. Expected Total Cost
Theorem (cf. J.Bertoin (1996) [1], Sato (1999) [8]).
$\tau_{b}^{+}=\inf\{t>0\mid X_{t}>b\}.$
Using this theorem, we will now examine the Laplace transforms of specific subordinators in detail.
@@ 3.1. Case where the Cumulative Demand is a Drifted Poisson Process
The Laplace transform of the FPT for a drifted Poisson process can be expressed using the Lambert W function.
@@
Under the assumptions of Theorem  3.4 , it holds that
(14) $E[T_{n}]$ $=\frac{a+(n-1)Q}{\mu+\alpha\lambda},$
(15) $\mathrm{Var}[T_{n}]$ $=\frac{\alpha^{2}\lambda(a+(n-1)Q)}{\left(\mu+\alpha\lambda\right)^{3}}.$
@@
be defined. We will organize the differentiation rules for the Laplace transform. We have
@@
We will also organize the second-order derivatives of the Laplace transform. We get
@@
Compute the first and second derivatives of $\Phi$. We have Let $B=\frac{\alpha}{\mu}$ and $C=\frac{\alpha\lambda}{\mu}$.
@@ 3.2. Case where the Cumulative Demand is a Drifted Compound Poisson Process
Recall the following result regarding the differentiation of the Laplace transform of $T_{n}$:
@@
$\frac{d^{2}\Phi}{ds^{2}}(s)$ $=-\frac{1}{2\mu}\times\frac{\left\{\sqrt{(\mu\eta+s+\lambda)^{2}-4\mu s\eta}-(s+\lambda-\mu\eta)\frac{(s+\lambda-\mu\eta)}{\sqrt{(\mu\eta+s+\lambda)^{2}-4\mu s\eta}}\right\}}{\left\{(\mu\eta+s+\lambda)^{2}-4\mu s\eta\right\}^{\frac{3}{2}}}$
@@
As a generalization of Theorem 3.7, we consider a compound Poisson process whose jump sizes follow a Gamma distribution. In this case, the Laplace transform cannot be expressed in closed form. However, since its derivatives at the origin can be computed explicitly, the expectation and variance can still be obtained.
@@
Let $L_{J_{1}}=\left(\frac{\eta}{\eta-\theta}\right)^{\beta}$. Then, for $\theta<\eta$, the Laplace Exponent $\Psi$ can be calculated as follows:
@@
Let $\psi^{-1}=\Phi$ denote the inverse function of $\psi$. Then, it holds that
$\Phi(0)=0\quad(\because\psi(0)=0),$
$\frac{d\Phi}{ds}(0)=\frac{1}{\cfrac{d\psi}{d\theta}(0)}=\frac{1}{\mu+\lambda\beta\eta^{-1}}=\frac{\eta}{\mu\eta+\beta\lambda},$
@@
Therefore, the expectation is given by
$E[T_{n}]$ $=-\left.\frac{d}{ds}E[e^{-sT_{n}}]\right|_{s=0}=K\exp(-K\Phi(0))=\frac{d\Phi}{ds}(0)\frac{\eta(a+(n-1)Q)}{\mu\eta+\beta\lambda}.$
The second moment is given by
@@ 3.3. Generalization of the Cumulative Demand
Let $\mu,\alpha,\lambda,\lambda^{\prime}>0$. Let $\{N_{t}\}_{t\geq 0}$ be a $\lambda$-Poisson process, $\{N^{\prime}_{t}\}_{t\geq 0}$ be a $\lambda^{\prime}$-Poisson process, and $\{J_{k}\}_{k\geq 1}$ be a sequence of independent and identically distributed random variables. Assume that $\{N_{t}\}_{t\geq 0}$, $\{N^{\prime}_{t}\}_{t\geq 0}$, and $\{J_{k}\}_{k\geq 1}$ are mutually independent. We define the cumulative demand process $\{D_{t}\}_{t\geq 0}$ as
@@
we derive $E[e^{-sT_{n}}]$, $E[T_{n}]$, and $\text{Var}[T_{n}]$. Let $\psi(\theta)$ denote the Laplace exponent of $D_{t}$. Then,
@@
Since $L_{J_{1}}=\left(\frac{\eta}{\eta-\theta}\right)^{\beta}$, for $\theta<\eta$, the Laplace Exponent $\Psi$ can be calculated as follows:
$\psi(\theta)=\log E[e^{\theta D_{1}}]=\theta\mu+\lambda(e^{\theta\alpha}-1)+\lambda^{\prime}\left(\left(\frac{\eta}{\eta-\theta}\right)^{\beta}-1\right),$
$\frac{d\psi}{d\theta}(\theta)=\mu+\alpha\lambda e^{\alpha\theta}+\lambda^{\prime}\beta\eta^{\beta}(\eta-\theta)^{-(\beta+1)},$
@@
Let $\Phi$ denote the inverse function of $\psi$. Then, it holds that
$\Phi(0)=0\quad(\because\psi(0)=0),$
$\frac{d\Phi}{ds}(0)=\frac{1}{\cfrac{d\psi}{d\theta}(0)}=\frac{1}{\mu+\alpha\lambda+\lambda^{\prime}\beta\eta^{-1}}=\frac{\eta}{\mu\eta+\alpha\lambda\eta+\beta\lambda^{\prime}},$
@@
Therefore, we have
$E[T_{n}]$ $=-\left.\frac{d}{ds}E[e^{-sT_{n}}]\right|_{s=0}=K\exp(-K\Phi(0))\frac{d\Phi}{ds}(0)=\frac{\eta(a+(n-1)Q)}{\mu\eta+\alpha\lambda+\beta\lambda^{\prime}},$
@@ 3.4. On the Expected Total Cost
Although we do not directly use the distribution of $T_{n}$, we consider the generalized cumulative demand process and investigate the expected value of the total inventory cost derived from it. Consider the cumulative demand process
@@
follows a Gamma distribution with parameters $j\beta,\eta$. Hence,
@@
Here, $\{N_{t}\}_{t\geq 0}$ is a $\lambda$-Poisson process, $\{N^{\prime}_{t}\}_{t\geq 0}$ is a $\lambda^{\prime}$-Poisson process, and $\{J_{k}\}_{k\geq 1}$ is a sequence of nonnegative independent and identically distributed random variables. Assume that these are mutually independent. Then, for any $s>0$ and $b>0$,
@@
holds.In particular, when $\beta=1$,
@@
Theorem 3.17 (Comparison between Finite-Horizon and Long-Run Average).
@@
Acknowledgements
The author is deeply grateful to Professor Seiji Hiraba and Takumu Ooi for their continuous support, insightful suggestions, and invaluable guidance. The author also thanks Professor Aya Ishigaki and Yuriko Ono for fruitful collaboration and stimulating discussions.